• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name brads

    Log entry time 13:25:30 on April12,2008

    Entry number 231610

    This entry is a followup to: 231609

    Followups:

    keyword=Chamber 2 U13 fixed itself

    We had to go back down for Bryan to check a trigger issue, so...

    I plugged in an A/D card and reconnected the signal cables between the card and the TDC. A quick cosmics run shows some reasonable data on U13 again...

    It seems to have fixed itself. Perhaps the field and signal wires shook themselves loose?


    FIGURE 1

    FIGURE 2